# Unforgettable Lessons from Forgettable Images: Intra-Class Memorability Matters in Computer Vision

### v4

@@ 1 Introduction
As we experience the world, our memory selectively retains certain objects while overlooking others [1, 2, 3]. For instance, among numerous flowers we encounter, a pale white daisy might fade from memory, while a striking violet osteospermum with intricate shapes remains vividly etched (Fig. 1). What factors make some flowers more memorable than others? To answer this question, we first introduce intra-class memorability which refers to the ability of instances within the same object class to be remembered. To delve into why certain items are more memorable than others, despite belonging to the same overall category, we conduct human behavior experiments using a continuous recognition task. In this task, participants are presented with a sequence of new and repeated images from the same class with delayed repetition and they are asked to identify those repeats. The proposed paradigm is novel compared to existing research, where continuous recognition tasks typically involve sequential presentations of images from different classes [4, 5, 6, 7, 8, 9].
To measure the memorability of individual instances, we introduce the Intra-Class Memorability score (ICMscore). Unlike prior evaluation metrics [7, 8, 10, 11], our ICMscore incorporates average hit and miss rates for an item’s repetitions, weighted by the time intervals between its initial and repeated presentations. Missing a repeat after a short interval indicates low memorability (LM), while remembering a repeat after a long interval signifies high memorability (HM).
Due to the lack of intra-class memorability datasets, we contribute the Intra-Class Memorability dataset (ICMD). The dataset contains 10 classes spanning over 500 images per class with their corresponding ICMscore aggregated from the responses of 2000 participants in the human behavior experiments on continuous recognition tasks. To demonstrate the effectiveness of our ICMD, we train and test AI models on different subsets of ICMD images based on their ICMscores across four downstream computer vision tasks: memorability prediction, image recognition, continual learning, and memorability-controlled image editing. Our experimental results show that a memorability prediction model trained on ICMD captures representative features diagnostic for intra-class memorability. Moreover, surprisingly, while highly memorable images seem more salient and thus easier to remember, we found that these images actually impair AI performance in image recognition and continual learning tasks. Finally, we use images from ICMD with and without masked semantic objects to fine-tune the state-of-the-art image diffusion model [12] using LoRA [13] to control image memorability. We highlight our key contributions:
@@ 2 Related Work
Research in cognitive psychology indicates that certain visual and contextual features—such as prominent visual characteristics, emotional valence, and semantic richness—substantially increase the likelihood of remembering specific objects over others [14]. This implies that memorability is a predictable and universal property, rather than purely subjective [15, 16, 17]. Based on this, Isola et al. [17] define “memorability" as the probability that a viewer will remember an image after a single viewing. Subsequent studies have investigated how intrinsic and extrinsic factors influence memorability. Extrinsic factors refer to an image’s categorical information, while intrinsic factors include other visual features beyond these categorical information [5, 6, 9]. However, none of these studies have effectively disentangled the effects of extrinsic and intrinsic features on memorability.
To address this problem, subsequent works [4, 7, 8, 18, 19] focus on specific super-categories, such as animals or indoor and outdoor scenes. In addition, some research has extended the studies on static images to videos, exploring memorability differences across video super-categories [20, 21, 22]. Although these studies limit categories to reduce the impact of extrinsic factors on memorability, they still mix images from different categories, making it difficult to conduct controlled studies that solely examine the effect of intrinsic features on memorability. To address this research gap, we introduce the concept of intra-class memorability. To investigate it, we propose a novel human behavior experimental paradigm focusing memorability among images within the same class.
Several memorability datasets have been collected through human behavior experiments [17, 4, 6, 7, 8], with LaMem [5] as one of the largest for deep learning-based image memorability prediction. In these studies, participants typically view a sequence of images, identifying any repeated target images. Memorability scores are often calculated based on the proportion of participants who correctly identify these repeats [17, 4, 6]. Some studies adjust scores to account for false alarms, where images are mistakenly identified as repeats [23, 7, 8]. The time interval between target presentations also influences memorability [5], yet random sampling of intervals can lead to unfair comparisons due to inconsistent controls. In our experiment, we standardize time intervals and introduce a novel memorability score, named as Intra-Class Memorability score (ICMscore). It accounts for both hit rates and miss rates of repeated images across a predefined set of time intervals. Moreover, we contribute a new dataset, Intra-Class Memorability dataset (ICMD). It consists of 5000 images with their corresponding ICMscores based on responses from over 2000 participants.
@@
Several studies have demonstrated that HM images can positively influence human cognitive states. For example, GANalyze [24] introduces a generative model that modifies image attributes to affect human emotional and cognitive responses. Peng et al. [25] show that memorable images boost user engagement in image retrieval systems and improve recall accuracy in storytelling and educational contexts. Similarly, [26] and [27] explore how AI-modified images and videos uniquely shape human memory. Thus, predicting memorability for a given image has become increasingly important [28, 29, 30]. However, existing models for predicting memorability rely on both extrinsic and intrinsic features, making them unable to capture fine-grained memorability differences within the same class of images. In contrast, our experiments demonstrate that our model, trained on the ICMD, outperforms existing predictors in accurately predicting intra-class memorability.
Despite early efforts to explore the role of memorability in enhancing human cognition, its impact on computer vision tasks remains largely unexplored. Specifically, the connection between image features relevant to object recognition and those that explain memorability is still unclear. Understanding this relationship could have significant implications for fields like continual learning. Using our ICMD, we present preliminary evidence that memorability affects AI performance in tasks such as image recognition and continual learning. To edit images with controls on memorability, we also build upon the existing generative models [12, 13] and introduce an image diffusion model, capable of manipulating image elements to enhance or reduce memorability.
@@ 3 Intra-Class Memorability Dataset(ICMD)
3.1 Human Behavior Experiments
Different from the existing memorability datasets [4, 5, 6, 9, 7, 8], our work introduces a new Intra-Class Memorability Dataset (ICMD), comprising naturalistic images from ten distinct object categories along with corresponding human memorability responses. The selected categories—volcano, dog, phone, lemon, car, bird, palace, flower, jellyfish, and teapot—were chosen to ensure clear semantic separation with no category overlap. We curated a total of 5,000 target images from ImageNet [31], evenly distributed across these ten classes. Fig. 2 presents the schematic of the human psychophysics experiments in the continuous object recognition tasks. In each trial, participants view a sequence of images, with both target and foil images belonging to the same category. Each image is followed by a blank screen to allow ample response time. Participants press the space bar when recognizing a previously seen image, with reaction times recorded during the 1,200 ms image presentation and the subsequent 1,600 ms blank screen. Before the start of an experiment, participants complete a guided demonstration to ensure task comprehension. During the experiment, immediate feedback follows each response: a green tick for correct answers and a red cross for incorrect ones. The image sequence is randomly shuffled, with target images reappearing after intervals of $t\in[8,16,24,32]$ images, uniformly distributed across these intervals. We determine these interval ranges through a pilot study exploring intervals $t$ from 2 to 64. The results reveal that intervals shorter than 8 led to near-ceiling performance, with participants achieving over 93% recall accuracy over all target images. Conversely, intervals exceeding 32 proved too challenging, with more than 34% of participants failing to identify any repeated images. Based on these findings, we select the interval range of [8, 16, 24, 32] to ensure meaningful image-dependent variability in memory recall performance. Each trial includes 100 unique target images, with foil images filling the sequence when targets are not shown, each target image and its repeat are shown only once to one participant. Sec. A.1 and Fig. A1 highlight differences in the experiment schematic between our study and previous works.
A total of 3,263 participants are recruited on Amazon Mechanical Turk (MTurk) [32]. To ensure data quality, participants are disqualified if they incorrectly identify more than three foil images as target images in a row or fail to respond within 60 seconds. After filtering the data from disqualified participants, we analyze responses from the remaining 2,000 qualified participants, resulting in 4,000 trials and a total of 346,537 key presses, with an average of 85 key presses per trial. The average false positive rate of the key presses is 15.24%. All experiments are conducted with informed consent and follow protocols approved by our institution’s Institutional Review Board.
@@ 3.2 Intra-Class Memorability Score (ICMscore)
To assess intra-class memorability, we introduce the Intra-Class Memorability Score (ICMscore), which measures the memorability of an image independently of its categorical information. Previous research typically computes memorability scores using hit rates and miss rates of memory recall [23, 7, 8]. However, the interval $t$ between image repetitions significantly affects memorability [5]. As shown in Fig. A2 and Sec. A.2, memory accuracy declines with longer intervals due to memory decays and increasing cognitive loads over time [33, 34]. Therefore, we incorporate the lengths of these time intervals into the ICMscore calculation.
@@
Since human memory retention diminishes over longer intervals, a longer interval $t$ makes correct identification of seen images more challenging. Therefore, when a participant correctly identifies an image after a longer interval, we assign it a higher memorability score $s$ to reflect the increased difficulty associated with the larger $t$. For correct responses ($x_{i}=1$), the memorability score $s_{i}$ is computed using a linearly increasing function: $s_{i}=x_{i}\frac{t_{i}}{T}$, where $t_{i}$ is the interval between two presentations of a target image, and $T=32$ is the maximum interval. Conversely, as longer time intervals are more prone to incorrect identifications of repeated target images, errors during shorter intervals are more indicative of the image’s memorability. For incorrect responses ($x_{i}=-1$), shorter intervals incur a larger penalty, while longer intervals result in a smaller penalty. The memorability score for incorrect responses is given by $s_{i}=x_{i}\frac{T-t_{i}}{T}$. The final ICMscore for each target image $I$ is calculated as the average of its memorability scores $s_{i}$ over all $n$ participants: ICMscore(I)$=\frac{\sum_{i=1}^{n}s_{i}}{n}$.
To highlight the importance of the temporal penalty in our ICMscore formulation, we conducted an additional analysis by removing the temporal component $t_{i}$. See Sec. A.5 for details. Results show that without this penalty, ICMscore fails to capture intra-class memorability effectively, diminishing its utility for downstream computer vision tasks. Furthermore, to demonstrate that ICMscore is not merely a variant of existing memorability metrics defined in [23, 7, 8], we applied an existing memorability prediction model trained on LaMem dataset [5] to estimate memorability scores for images in our ICMD dataset. The two sets of scores show only a weak positive correlation, with a Pearson coefficient of 0.26. This suggests that ICMscore captures distinct memorability-related features overlooked by prior metrics and serves a complementary role in memorability evaluation.
Figure 3: Correlation analysis and Grad-CAM visualization. (a) Pearson correlation between feature distances of images to their categorical centroids and their ICMscore. The y-axis shows correlation values while the x-axis displays class labels. Correlations range from $r=0.21$ for “Flower" to $r=0.21$ for “Phone" ($p<0.05$ for all ten classes). (b) Grad-CAM visualizations displaying importance maps for three high memorability (HM) images. These visualizations highlight key regions by computing gradients of the memorability predictor with respect to feature maps to identify areas contributing most to memorability scores. For HM images, features enhancing memorability are highlighted according to the color scale shown in the right color bar. The first column presents original images, the second column shows visualizations from the model trained on LaMem dataset [5], and the third column displays results from our model trained on ICMD.
@@ 3.3 Dataset Analysis
Given the target images and their corresponding ICMscores, we now analyze how visual features of target images influence ICMscores. To investigate this relationship, we use a ResNet [35] model pre-trained on ImageNet [31] for object recognition tasks to extract feature representations for each image. We define the mean of all image features belonging to the same category as the centroid. Next, we compute the Pearson correlation [36] between the L2 distances from each centroid to individual target image features and their ICMscores.
@@ 4.1 Memorability Prediction
Memorability prediction can provide insights into human cognition and perception by helping us understand how and why certain images are more memorable than others [37, 38, 39]. Existing memorability prediction methods fail to disentangle the intrinsic and extrinsic factors that influence the memorability of an image. In contrast, our work focuses on predicting ICMscore based exclusively on intrinsic factors of images within the same category.
@@
Specifically, we use a ResNet [35], pre-trained on ImageNet [31] for image recognition tasks, to extract image features, followed by two fully connected layers to predict memorability scores. We train our ICMscore predictor using the ICMD with an L2 regression loss, where the ICMscore of the input image serves as the ground truth. To validate the effectiveness of our predictor, we compare it against two baseline models. Baseline 1 employs the same network structure as our model but is trained from scratch without loading the pre-trained network parameters. For Baseline 2, inspired by previous work [30, 40], we use a Vision Transformer (ViT) pretrained on ImageNet [41], append two randomly initialized fully connected layers, and fine-tune only these added layers while keeping the ViT backbone frozen. We perform 5-fold cross-validation to ensure the statistical robustness of the results. Following the practices in memorability prediction [10, 30, 5], we assess the models’ performance by measuring the Spearman rank correlation [42] between the predicted and ground truth ICMscores. The first model achieves the highest correlation (0.64 ± 0.09), demonstrating that ICMscores can be effectively predicted using features extracted from pre-trained image recognition models, while the other two models obtained lower correlations: 0.60±0.2 and 0.57±0.05.
@@
To ensure the 0.64 correlation from our predictor is not due to model overfitting, we split human participants into two independent groups and computed the Spearman correlation of their ICMscores on the same target images. Across 10 runs, the average correlation is $0.69\pm 0.05$, demonstrating high consistency in within-human memorability. Although our model’s correlation is slightly lower than human consistency, it exceeds both baselines, suggesting that the model effectively captures diagnostic visual features of memorability rather than overfitting the data. This finding also aligns with the observed positive correlation between feature distance from the categorical centroids and ICMscores (Sec 3.3).
@@
To further explore what image components contribute to high memorability, we employ Grad-CAM [43] to visualize importance maps for memorability, as shown in Fig. 3(b). For highly memorable (HM) images, our predictor focuses on distinctive regions. For instance, in a teapot image with an ICMscore of 0.78, our predictor emphasizes the unique face-like texture of the teapot, while the model trained on LaMem [5] focuses on the teapot’s spout and handle which are typical features reflecting the generic categorical appearance of teapots.

### v5

@@ 1 Introduction
As we experience the world, our memory selectively retains certain objects while overlooking others McDermott & Roediger (2018); Halamish & Stern (2022); Leshikar et al. (2012). For instance, among numerous flowers we encounter, a pale white daisy might fade from memory, while a striking violet osteospermum with intricate shapes remains vividly etched (Fig. 1). What factors make some flowers more memorable than others? To answer this question, we first introduce intra-class memorability which refers to the ability of instances within the same object class to be remembered. To delve into why certain items are more memorable than others, despite belonging to the same overall category, we conduct human behavior experiments using a continuous recognition task. In this task, participants are presented with a sequence of new and repeated images from the same class with delayed repetition and they are asked to identify those repeats. The proposed paradigm is novel compared to existing research, where continuous recognition tasks typically involve sequential presentations of images from different classes Bylinskii et al. (2015); Khosla et al. (2015); Dubey et al. (2015); Goetschalckx & Wagemans (2019); Kramer et al. (2023); Hovhannisyan et al. (2021).
To measure the memorability of individual instances, we introduce the Intra-Class Memorability score (ICMscore). Unlike prior evaluation metrics Goetschalckx & Wagemans (2019); Kramer et al. (2023); Isola et al. (2014); Akagunduz et al. (2019), our ICMscore incorporates average hit and miss rates for an item’s repetitions, weighted by the time intervals between its initial and repeated presentations. Missing a repeat after a short interval indicates low memorability (LM), while remembering a repeat after a long interval signifies high memorability (HM).
@@
Due to the lack of intra-class memorability datasets, we contribute the Intra-Class Memorability dataset (ICMD). The dataset contains 10 classes spanning over 500 images per class with their corresponding ICMscore aggregated from the responses of 2000 participants in the human behavior experiments on continuous recognition tasks. To demonstrate the effectiveness of our ICMD, we train and test AI models on different subsets of ICMD images based on their ICMscores across four downstream computer vision tasks: memorability prediction, image recognition, continual learning, and memorability-controlled image editing. Our experimental results show that a memorability prediction model trained on ICMD captures representative features diagnostic for intra-class memorability. Moreover, surprisingly, while highly memorable images seem more salient and thus easier to remember, we found that these images actually impair AI performance in image recognition and continual learning tasks. Finally, we use images from ICMD with and without masked semantic objects to fine-tune the state-of-the-art image diffusion model Brooks et al. (2022) using LoRA Hu et al. (2021) to control image memorability. We highlight our key contributions:
@@ 2 Related Work
Research in cognitive psychology shows that certain visual and contextual features, such as salient image characteristics, emotional valence, and semantic richness, significantly increase the likelihood of remembering specific objects Schmidt (1991). This suggests that memorability is a predictable and universal property rather than a purely subjective one Konkle et al. (2010); Bylinskii et al. (2022); Isola et al. (2011). Building on this, Isola et al. Isola et al. (2011) define “memorability” as the probability that an image will be remembered after a single viewing. Subsequent studies have examined both extrinsic factors, such as category information, and intrinsic factors, such as visual features beyond category Khosla et al. (2015); Dubey et al. (2015); Hovhannisyan et al. (2021). Yet, these works have not disentangled the respective contributions of extrinsic and intrinsic factors.
To mitigate extrinsic influences, later studies restricted their focus to specific super-categories such as animals or indoor versus outdoor scenes Bylinskii et al. (2015); Goetschalckx & Wagemans (2019); Kramer et al. (2023); Lu et al. (2020); Borkin et al. (2013), or extended memorability analysis to videos across super-categories Cohendet et al. (2018a); Newman et al. (2020); Cohendet et al. (2018b). However, these approaches still mix images from different categories, limiting control over intrinsic features. To address this gap, we introduce intra-class memorability, which examines why some images are more memorable than others within the same category. We design a novel human behavioral paradigm specifically targeting this dimension.
Prior memorability datasets Isola et al. (2011); Bylinskii et al. (2015); Dubey et al. (2015); Goetschalckx & Wagemans (2019); Kramer et al. (2023), with LaMem Khosla et al. (2015) as one of the largest for deep learning-based memorability prediction, have relied on repeated-image recognition tasks, with scores typically derived from the proportion of correct detections and sometimes adjusted for false alarms Bainbridge & Rissman (2018); Goetschalckx & Wagemans (2019); Kramer et al. (2023). While temporal spacing between repeats is known to affect memorability Khosla et al. (2015), prior protocols often used random intervals, introducing uncontrolled variability. In contrast, our experiment standardizes time intervals and introduces the Intra-Class Memorability score (ICMscore), which jointly accounts for hit and miss rates across predefined intervals. We further contribute the Intra-Class Memorability Dataset (ICMD), comprising 5,000 images with ICMscores derived from responses of over 2,000 participants.
@@
Studies have shown that highly memorable (HM) images can positively influence human cognition. For instance, GANalyze Goetschalckx et al. (2019) uses generative models to alter image attributes and modulate emotional or cognitive responses. Peng et al. Peng & Bainbridge (2024) demonstrate that memorable images increase engagement in retrieval systems and improve recall in storytelling and education. Other works Pataranutaporn et al. (2024); Siarohin et al. (2019) further reveal how AI-modified images and videos shape memory. These findings highlight the growing importance of memorability prediction Lahrache & El Ouazzani (2022); Squalli-Houssaini et al. (2018); Hagen & Espeseth (2023). However, existing predictors depend on both extrinsic and intrinsic features and cannot capture fine-grained differences within a class. In contrast, our model, trained on ICMD, achieves more accurate intra-class memorability prediction.
While memorability has been studied in human cognition, its impact on computer vision tasks remains underexplored. In particular, the link between image features critical for recognition and those that explain memorability is unclear, with implications for continual learning. Using ICMD, we provide initial evidence that memorability influences AI performance in recognition and continual learning. Finally, to enable direct manipulation, we fine-tune the pre-trained InstructPix2Pix model (Ip2p) Brooks et al. (2022) with LoRA Hu et al. (2021), enabling controlled manipulation of the memorability of specific object classes.
@@ 3 Intra-Class Memorability Dataset(ICMD)
3.1 Human Behavior Experiments
In contrast to prior memorability datasets Bylinskii et al. (2015); Khosla et al. (2015); Dubey et al. (2015); Hovhannisyan et al. (2021); Goetschalckx & Wagemans (2019); Kramer et al. (2023), we introduce the Intra-Class Memorability Dataset (ICMD), which pairs naturalistic images with human memorability annotations across ten object categories. The categories—volcano, dog, phone, lemon, car, bird, palace, flower, jellyfish, and teapot—were selected to ensure clear semantic separation without overlap. ICMD comprises 5,000 target images curated from ImageNet Deng et al. (2009), evenly distributed across these classes.
@@ 3.1 Human Behavior Experiments
Fig. 2 illustrates the schematic of our human psychophysics experiments. In each trial, participants viewed a sequence of images, with both targets and foils drawn from the same category. Each image was displayed for 1,200 ms followed by a 1,600 ms blank screen, during which participants could press the spacebar to indicate recognition of a previously seen image. Reaction times were recorded across both periods. Before the main experiment, participants completed a guided demonstration to ensure task comprehension. During the experiment, immediate feedback was provided after each response (green tick for correct, red cross for incorrect). Target images reappeared after intervals of $t\in{8,16,24,32}$ images, sampled uniformly. These values were chosen based on a pilot study testing intervals from 2 to 64: intervals shorter than 8 led to near-ceiling performance (over 93% recall), while intervals longer than 32 were too difficult, with more than 34% of participants failing to detect any repeats. Each trial contained 100 unique targets, with foils filling the remaining sequence; each target and its repeat were shown only once per participant. Sec. A and Fig. A1 further highlight differences between our setup and prior studies.
We recruited 3,263 participants on Amazon Mechanical Turk (MTurk) Turk (2012). To ensure data quality, participants were disqualified if they misidentified more than three consecutive foils as targets in a row or failed to respond within 60 seconds. After filtering, responses from 2,000 qualified participants remained, yielding 4,000 trials and 346,537 key presses, with an average of 85 key presses per trial. The mean false positive rate was 15.24%. All experiments were conducted with informed consent under protocols approved by our Institutional Review Board.
@@ 3.2 Intra-Class Memorability Score (ICMscore)
To assess intra-class memorability, we introduce the Intra-Class Memorability Score (ICMscore), which measures the memorability of an image independently of its categorical information. Previous research typically computes memorability scores using hit rates and miss rates of memory recall Bainbridge & Rissman (2018); Goetschalckx & Wagemans (2019); Kramer et al. (2023). However, the interval $t$ between image repetitions significantly affects memorability Khosla et al. (2015). As shown in Fig. A2 and Sec. B, memory accuracy declines with longer intervals due to memory decays and increasing cognitive loads over time Wixted (2004); Sikarwar & Zhang (2024). Therefore, we incorporate the lengths of these time intervals into the ICMscore calculation.
@@
[…]
To highlight the importance of the temporal penalty in our ICMscore formulation, we conducted an additional analysis by removing the temporal component $t_{i}$. Results show that without this penalty, ICMscore fails to capture intra-class memorability effectively, diminishing its utility for downstream computer vision tasks. See Sec. C and Fig. A3 for details. Furthermore, to demonstrate that ICMscore is not merely a variant of existing memorability metrics defined in Bainbridge & Rissman (2018); Goetschalckx & Wagemans (2019); Kramer et al. (2023), we applied an existing memorability prediction model trained on LaMem dataset Khosla et al. (2015) to estimate memorability scores for images in our ICMD dataset. The two sets of scores show only a weak positive correlation, with a Pearson coefficient of 0.26. This suggests that ICMscore captures distinct memorability-related features overlooked by prior metrics and serves a complementary role in memorability evaluation.
Figure 3: Correlation analysis and Grad-CAM visualization. (a) Pearson correlation between image feature distances to category centroids and their ICMscores. The y-axis indicates correlation values and the x-axis shows class labels, ranging from $r=0.21$ for “flower” to $r=0.21$ for “phone” ($p<0.05$ for all classes). (b) Grad-CAM visualizations of three high-memorability (HM) images. Importance maps are derived from gradients of the memorability predictor with respect to feature maps, highlighting regions that contribute most to memorability scores. For HM images, memorability-enhancing features are highlighted according to the color bar on the right. The first row shows original images, the second row shows results from the LaMem-trained model Khosla et al. (2015), and the third row shows results from our ICMD-trained model.
@@ 3.3 Dataset Analysis
Given the target images and their corresponding ICMscores, we now analyze how visual features of target images influence ICMscores. To investigate this relationship, we use a ResNet He et al. (2015) model pre-trained on ImageNet Deng et al. (2009) for object recognition tasks to extract feature representations for each image. We define the mean of all image features belonging to the same category as the centroid. Next, we compute the Pearson correlation Sedgwick (2012) between the L2 distances from each centroid to individual target image features and their ICMscores.
@@ 4.1 Memorability Prediction
Memorability prediction offers insights into human cognition and perception by revealing why some images are remembered more easily than others Perera et al. (2019); Harada & Sakai (2023); Needell & Bainbridge (2022). However, existing approaches conflate intrinsic and extrinsic factors. In contrast, our work predicts ICMscore by isolating intrinsic factors within the same category.
@@
Specifically, we use a ResNet He et al. (2015), pre-trained on ImageNet Deng et al. (2009) for image recognition tasks, to extract image features, followed by two fully connected layers to predict memorability scores. We train our ICMscore predictor using the ICMD with an L2 regression loss, where the ICMscore of the input image serves as the ground truth. To validate the effectiveness of our predictor, we compare it against two baseline models. Baseline 1 employs the same network structure as our model but is trained from scratch without loading the pre-trained network parameters. For Baseline 2, inspired by previous work Hagen & Espeseth (2023); Constantin & Ionescu (2021), we use a Vision Transformer (ViT) pretrained on ImageNet Alexey (2020), append two randomly initialized fully connected layers, and fine-tune only these added layers while keeping the ViT backbone frozen. We perform 5-fold cross-validation to ensure the statistical robustness of the results. Following the practices in memorability prediction Isola et al. (2014); Hagen & Espeseth (2023); Khosla et al. (2015), we assess the models’ performance by measuring the Spearman rank correlation Zar (2005) between the predicted and ground truth ICMscores. The first model achieves the highest correlation (0.64 ± 0.09), demonstrating that ICMscores can be effectively predicted using features extracted from pre-trained image recognition models, while the other two models obtained lower correlations: 0.60±0.2 and 0.57±0.05.
@@
To verify that the 0.64 correlation from our predictor is not due to overfitting, we split participants into two independent groups and computed the Spearman correlation of their ICMscores on the same images. Averaged over 10 runs, the correlation was $0.69\pm 0.05$, indicating high within-human consistency. While our model falls slightly below this human upper bound, it outperforms both baselines, suggesting it captures diagnostic visual features of memorability rather than overfitting. This result also aligns with the positive correlation between feature distance from category centroids and ICMscores (Sec. 3.3).
@@
To further explore what image components contribute to high memorability, we employ Grad-CAM Selvaraju et al. (2017) to visualize importance maps for memorability, as shown in Fig. 3(b). For highly memorable (HM) images, our predictor focuses on distinctive regions. For instance, in a teapot image with an ICMscore of 0.78, our predictor emphasizes the unique face-like texture of the teapot, while the model trained on LaMem Khosla et al. (2015) focuses on the teapot’s spout and handle which are typical features reflecting the generic categorical appearance of teapots.